Klean Fiber Company is the creator of Y-Go, a technology that weaves silver into its fabrics to kill bacteria and odor on clothing while managing heat. Y-Go has become very popular in undergarments for sports activities. Operating at capacity, the company can produce 1,046,000 Y-Go undergarments a year. The per unit and the total costs for an individual garment when the company operates at full capacity are as follows.

Per Undergarment Total
Direct materials $1.99 $2,081,540
Direct labor 0.48 502,080
Variable manufacturing overhead 1.03 1,077,380
Fixed manufacturing overhead 1.53 1,600,380
Variable selling expenses 0.30 313,800
   Totals $5.33

$5,575,180

The U.S. Army has approached Klean Fiber and expressed an interest in purchasing 249,900 Y-Go undergarments for soldiers in extremely warm climates. The Army would pay the unit cost for direct materials, direct labor, and variable manufacturing overhead costs. In addition, the Army has agreed to pay an additional $0.97 per undergarment to cover all other costs and provide a profit. Presently, Klean Fiber is operating at 70% capacity and does not have any other potential buyers for Y-Go. If Klean Fiber accepts the Army’s offer, it will not incur any variable selling expenses related to this order.


Prepare an incremental analysis for the Klean Fiber. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Reject
Order
Accept
Order
Net Income
Increase
(Decrease)
Revenues $

$

$

Variable costs:
   Direct materials

   Direct labor

   Variable overhead

   Total variable costs

Net income $

$

$



Should Klean Fiber accept the Army's offer?

Klean Fiber should

accept/reject

the Army’s offer.
Answer #1
Reject order Accept order Net income increase(decrease)
Revenues 0 1117053 1117053
Variable costs:
Direct materials 0 497301 -497301
Direct labor 0 119952 -119952
Variable overhead 0 257397 -257397
Total variable costs 0 874650 -874650
Net income 0 242403 242403
Klean Fiber should accept the Army’s offer.
Formuals used:
Revenues 1117053 =249900*(1.99+0.48+1.03+0.97)
Variable costs:
Direct materials 497301 =249900*1.99
Direct labor 119952 =249900*0.48
Variable overhead 257397 =249900*1.03